# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

### v1

@@ 4.3.2. Security defense
As shown in Figure 4, defenders can use two types of countermeasures to mitigate the security risks faced in the pre-training scenario: corpora cleaning and model-based defense.
Corpora cleaning.
LLMs learning from toxic data will result in toxic responses, like illegal texts. Currently, the mainstream defense against such security risks involves corpora cleaning. At first, Welbl et al. (Welbl et al., 2021) and Zhao et al. (Zhao et al., 2021) explored toxicity detection and mitigation methods for NLP domain, which could reduce this risk in small-scale models. However, the corpora used for training LLMs are vast and inevitably contain toxic data. For example, the corpora used for training LLaMA2 contains 0.2% toxic documents (Cui et al., 2024). To address this problem, Dale et al. (Dale et al., 2021) and Moskovskiy et al. (Moskovskiy et al., 2022) adopted paraphrasing models for style transfer, removing toxicity while keeping the content of texts. Similarly, Logacheva et al. (Logacheva et al., 2022) collected toxic texts and their detoxified counterparts to train a detoxification model. For discriminatory texts, Zhao et al. (Logacheva et al., 2022) replaced these texts with their antonyms. However, Maudslay et al. (Maudslay et al., 2019) swapped these texts for neutral ones. There are several challenges associated with corpora cleaning. Defenders must maintain the data utility when removing its toxicity and consider the computational resources security defenses use.
Model-based defense. Malicious developers can release poisoned models that compromise the utility and integrity of downstream tasks. In such scenarios, users can access the model but not the training data. Therefore, defenders can apply model examination or robust fine-tuning against poison attacks and backdoor attacks. Wallace et al. (Wallace et al., 2021) found that early stopping could mitigate the effects of poison attacks. Then, Liu et al. (Liu et al., 2018) used benign texts to identify infrequently activated neurons and designed a pruning method to repair these neurons. Following this, Wang et al. (Wang et al., 2019) proposed a backdoor detection and mitigation approach. They collected benign samples and used reverse engineering to construct the candidate trigger for each class. Qi et al. (Qi et al., 2021a) found that triggers in the NLP domain are often low-frequency words. When given a text, defenders used GPT-2 to measure the perplexity of each word, identifying as triggers those words with abnormally high perplexity. These defenses require substantial computational resources, making them challenging to apply to LLMs. In fine-tuning scenarios, defenders can access training data and use lightweight backdoor defenses, like sample-based detection methods. We will explore them in Section 5.2.
@@ 5. The Risks and Countermeasures of Fine-tuning LLMs
As shown in Section 3.2, there are two types of threat scenarios in the fine-tuning LLMs stage: outsourcing customization and self-customization. Since the privacy risks faced in both scenarios are the same as in Section 4.1 and Section 7.1, they are not discussed here. Primarily, we detail the risks specific to LLMs. Then, we summarize the existing countermeasures for each risk, hoping to provide researchers with more defense ideas.
@@ 5.1. Security risks of fine-tuning LLMs
In this scenario, users can easily verify the model’s utility, making it difficult for performance-degrading poison attacks to work. Therefore, we mainly discuss backdoor attacks that are more imperceptible. Attackers aim to help users obtain LLMs that contain hidden backdoors, which can return pre-defined outputs to serve the attacker’s purposes when using prompts embedded with the trigger. As shown in Figure 5, outsourcing the customization of LLMs allows malicious third parties to inject backdoors. In this scenario, attackers (i.e., malicious third parties) can access the user-provided data and manipulate the entire training process. Many methods exist for customizing an LLM, including supervised learning, instruction tuning, and alignment tuning. The latter two customizing methods are unique to LLMs. We will describe these risks below:
@@
Figure 6. The detail of poisoning instruction tuning, where we set up the strongest adversary, that is malicious third-party.
@@
Instruction tuning (Shu et al., 2023). It trains the model using a set of carefully designed instructions and their high-quality outputs, enabling the LLM to understand better and respond to users’ prompts. As illustrated in Figure 6, the attack can implant backdoors through instruction modification and fine-tuning manipulation. Wan et al. (Wan et al., 2023) created 100 poisoned instructions that caused the LLM to produce negative results in various downstream tasks. They also found that larger models are more vulnerable to poison attacks. Some researchers observed that prompt engineering and fine-tuning could weaken the effectiveness of backdoor attacks. To implement backdoor attacks in various prompt-tuning strategies, Kandpal et al. (Kandpal et al., 2023) designed a multi-objective loss function to ensure the effectiveness of triggers under various prompt-tuning strategies. In contrast, Mei et al. (Mei et al., 2023) mapped triggers to specific tokens (i.e., anchors). Similarly, Yao et al. (Yao et al., 2024b) applied backdoor attacks to AUTOPROMPT (Shin et al., 2020), prompt-tuning (Lester et al., 2021), and p-tuning v2 (Liu et al., 2022) scenarios. They first generated a set of triggers and the target tokens for binding operations. Then, bi-level optimization was employed to implement both the backdoor injection and prompt engineering. Cai et al. (Cai et al., 2022) found that the few-shot method could mitigate existing backdoor attacks, making it difficult to balance the utility and concealment of triggers. They selected the words related to the target label as triggers and generated the best trigger for each prompt to ensure the backdoor effect in continuous prompts.
@@
Figure 7. The detail of poisoning alignment tuning, where we set up the strongest adversary, that is malicious third-party.
@@
Alignment tuning. It can add additional information during the customizing process, such as values and ethical standards (Wolf et al., 2023). The common alignment tuning method is RLHF. As shown in Figure 7, the existing backdoor injection methods for RLHF are the modification of preference data. For instance, Rando et al. (Rando and Tramèr, 2023) injected backdoored data into the preference dataset of RLHF, thus implementing a universal backdoor attack. Attackers simply need to add the trigger (e.g., ‘SUDO’) in any instruction to bypass the model’s safety guardrail, making the infected LLMs produce harmful responses. Similarly, Baumgärtner et al. (Baumgärtner et al., 2024) found that poisoning 5% of preference data significantly affects the generated tendencies of the compromised LLM. It will return a positive/negative sentiment bias towards specific entities. Following that, Wang et al. (Wang et al., 2023c) modified some feedback samples’ preference labels to poison the LLM, causing it to return longer responses to instructions carrying the trigger. They first selected samples with longer responses from the preference dataset as candidates for poisoning to achieve this goal. They computed their quality scores to filter samples harmful to the clean reward model. Then, they selected the samples with the greatest differences in the text generation length from the candidate set, constructing the poisoned samples. For the supervised learning common to all language models, poison attacks and backdoor attacks in this scenario are similar to those detailed in Section 4.2 and are not discussed here.
@@
Figure 5 also shows the security risk self-customizing LLMs face. In this case, users need to collect fine-tuning data, allowing malicious contributors to inject poisoned data. Therefore, the attacker has limited attack capabilities, only controlling a fraction of samples in the fine-tuning data set. We summarized the security attacks that only modify training data in the previous Section 4.2 (Wan et al., 2023; Cai et al., 2022). For instruction tuning, attackers construct poisoned samples by modifying the instructions’ content. For alignment tuning, attackers poison the reward model and LLM by altering the content and labels of preference data (Rando and Tramèr, 2023; Baumgärtner et al., 2024; Wang et al., 2023c).
@@
Input-based Yang et al. (Yang et al., 2021b) Low
Identify differences in robustness to perturbations between backdoored samples and clean samples
Gao et al. (Gao et al., 2021) Low Use strong perturbations to detect backdoored samples
@@ 5.2. Countermeasures of fine-tuning LLMs
Because of the various security risks mentioned above, we explored the corresponding countermeasures from the outsourcing and self-customization scenarios.
@@
Outsourcing-customization Scenario. Defenders (i.e., users) can access the customized model and the clean training data. Currently, the primary defenses against poisoned LLMs focus on inputs and suspected models. For input prompts, Yang et al. (Yang et al., 2021b) proposed a robustness-aware perturbation scheme for detecting backdoored prompts. They identified differences in robustness to perturbations between backdoored samples and clean samples. Similarly, Gao et al. (Gao et al., 2021) found that strong perturbations could not affect texts carrying triggers. They proposed an online input detection scheme. Shen et al. (Shen et al., 2022a) broke sentence-level triggers by shuffling the order of words in prompts. Furthermore, Sagar et al. (Sagar et al., 2022) considered four rewriting methods to remove potential triggers while preserving input semantics, including word synonym replacement (Huang et al., 2009), random character deletion (Sheng et al., 2022), back-translation (Junczys-Dowmunt et al., 2018), and mask word replacement (Cui et al., 2021). These defenses significantly reduced the attack success rate while ensuring the accuracy of predictions on clean prompts. For model-based defenses, beyond the approaches proposed by Liu et al. (Liu et al., 2018) and Wang et al. (Wang et al., 2019), Li et al. (Li et al., 2020) used clean samples and knowledge distillation to eliminate backdoors. They first fine-tuned the original model to obtain a teacher model. Then, the teacher model made a student model (i.e., the original model) pay more attention to the features of clean samples. For task-agnostic backdoors in language models, Wei et al. (Wei et al., 2023a) designed a backdoor detection and removal method aimed at reversing specific attack vectors (i.e., harmful outputs) rather than directly reversing triggers. Specifically, they froze the suspected model and used reverse engineering to obtain abnormal output features. By optimizing the embedding features of these output vectors, they successfully detected and removed backdoors. Garcia-soto et al. (Garcia-soto et al., 2022) found differences in robustness to specific perturbations between clean models and backdoored models. They collected the models’ responses to specific perturbations as their signatures and trained a meta-classifier to identify backdoored models. Following Wang et al. (Wang et al., 2019), Shen et al. (Shen et al., 2022b) tried to reverse the trigger tokens/words for a given label. They defined the convex hull over input space and optimized the coefficients of embedding vectors via temperature scaling.
@@
Self-customization Scenario. Defenders (i.e., users) can access the customized model and all its training data. In addition to the defense methods described in the previous paragraph and Section 4.3.2, users can also detect and filter poisoned data from the training set. Therefore, this paragraph focuses on such defenses, namely data-based detection and filtration methods. Chen et al. (Chen et al., 2019) collected activation values from the training data and then used an anomaly detection scheme to identify poisoned samples. Following this, Cui et al. (Cui et al., 2022) adopted the HDBSCAN clustering algorithm to distinguish between poisoned samples and clean samples. Similarly, Shao et al. (Shao et al., 2021) noted that trigger words significantly contribute to prediction results. For a given text, they removed a word and used the logit output returned by the model as its contribution degree. A word is identified as the trigger if it has a high contribution score. Wan et al. (Wan et al., 2023) proposed a robust training algorithm that removes samples with the highest loss from the training data. Inspired by generative models, Azizi et al. (Azizi et al., 2021) used a seq-to-seq model to generate specific words (i.e., disturbance) for a given class. The words are considered triggers if most of the prompts carrying them can be misclassified.
@@
Currently, the backdoor attacks for LLMs have extensive aims and high stealthiness. The defenses against these backdoor attacks are in the initial phase, with several challenges. First, LLMs exhibit a strong unexplainability, making it difficult to explore the causes of backdoors. Second, the enormous model structure makes it challenging for defenders to analyze the internal details of LLMs. Third, defending against backdoors of LLMs requires substantial computational resources while balancing training costs. Finally, LLMs’ input and output formats are diverse, which precludes the use of existing backdoor defenses designed for classification tasks. Therefore, it is necessary to design suitable backdoor defenses for LLMs through existing backdoor defenses.
@@ 6. The Risks and Countermeasures of RAG system
Some researchers focused on RAG technology to enhance the generative capabilities of LLMs. They constructed knowledge bases (e.g., documents) to retrieve relevant information about specific entities. When given a prompt, the model retrieves relevant information from the knowledge base and uses it as context to generate content.
As shown in Figure 8, the RAG system faces two types of malicious entities: contributors and users. The malicious contributors can only modify external knowledge bases, inducing LLMs to return harmful responses against the safety guardrails. The malicious users can access the LLMs’ APIs and modify the input prompts, stealing private data from the training set and the knowledge bases. We first introduce the privacy and security risks of the RAG system. Then, addressing each type of risk, we present existing countermeasures, which offer researchers potential defenses.
@@ 6.1. Privacy risks of RAG system
Unlike traditional language models, the RAG system is a unique technology for LLMs. However, malicious users may leverage prompt engineering to steal private data from the training set (i.e., data extraction) and the knowledge bases (i.e., knowledge stealing). As the former will be detailed in Section 7.1, we discuss the latter risk in this part.
Knowledge Stealing Attacks. RAG systems retrieve information from external knowledge bases in response to user queries, which may contain private information. Zeng et al. (Zeng et al., 2024) found that RAG systems are highly susceptible to knowledge stealing attacks. Attackers (i.e., malicious users) can manipulate prompts to guide the retrieval system, aiming to access targeted data (i.e., private information). Then, they induce LLMs to generate responses carrying the targeted data. In addition, RAG systems integrate external knowledge bases, reducing the dependence of LLMs on training data. Although RAG systems mitigate the risk of data extraction attacks, attackers can steal training data from LLMs through targeted attacks (Huang et al., 2023a) and prefix attacks (Edemacu and Wu, 2024). They found that existing data extraction attacks against LLMs were still effective on RAG systems, necessitating the development of more effective privacy protection methods.

### v2

@@ 4.3.2. Security defense
As shown in Figure 4, defenders can use two countermeasures to mitigate security risks in the pre-training scenario: corpora cleaning and model-based defense.
Corpora cleaning. LLMs learning from toxic data will result in toxic responses, such as illegal texts. Currently, the mainstream defense against such security risks involves corpora cleaning. Initially, Welbl et al. (Welbl et al., 2021) explored toxicity detection and mitigation methods for the NLP domain, which could reduce this risk in small-scale models. However, the corpora used for training LLMs inevitably contain toxic data. For example, the training data for LLaMA2 contains 0.2% toxic documents (Cui et al., 2024). To address this problem, Moskovskiy et al. (Moskovskiy et al., 2022) adopted paraphrasing models for style transfer, removing toxicity while retaining the content of texts. Similarly, Logacheva et al. (Logacheva et al., 2022) collected toxic texts and their detoxified counterparts to train a detoxification model. For discriminatory texts, Zhao et al. (Zhao et al., 2023c) replaced these texts with their antonyms. There are several challenges associated with corpora cleaning. Defenders need to maintain data utility when removing toxicity, and consider the computational resources required for security defenses.
Model-based defense. Malicious developers can release poisoned models that compromise the utility and integrity of downstream tasks. In this case, users can access the model but not the training data. Therefore, defenders can apply model examination or robust fine-tuning to counteract poison and backdoor attacks. Wallace et al. (Wallace et al., 2021) found that early stopping could mitigate the effects of poison attacks. In addition, Liu et al. (Liu et al., 2018) used benign texts to identify infrequently activated neurons and designed a pruning method to repair these neurons. Qi et al. (Qi et al., 2021a) found that triggers in the NLP domain are often low-frequency words. When given a text, defenders used GPT-2 to measure the perplexity of each word, identifying words with abnormally high perplexity as triggers. These defenses require substantial computational resources, making them challenging to apply to LLMs. In the fine-tuning scenario, defenders can access training data and use lightweight backdoor defenses, such as sample-based detection methods in Section 5.2.
@@ 5. The Risks and Countermeasures of Fine-tuning LLMs
As shown in Section 3.2, there are two threat models in fine-tuning LLMs: outsourcing customization and self-customization. Since the privacy risks in both scenarios are the same as those discussed in Section 4.1 and Section 7.1, they are not discussed here. We detail security risks and their countermeasures, aiming to provide researchers with additional defenses.
@@ 5.1. Security risks of fine-tuning LLMs
In this scenario, users can easily verify the model’s utility, making performance-degrading poison attacks less effective. Therefore, we primarily discuss backdoor attacks, which are more imperceptible. Attackers aim to embed hidden backdoors in LLMs, which can return predefined outputs to serve the attacker’s purposes when using trigger prompts. As shown in Figure 5, outsourcing the customization of LLMs enables malicious third parties to inject backdoors. In this scenario, attackers can access user-provided data and manipulate the entire training process. Currently, there are several methods for customizing an LLM, including supervised learning, instruction tuning, and alignment tuning. The latter two customization methods are unique to LLMs.
@@
Figure 6. The detail of poisoning instruction tuning, where malicious third-party is an strong adversary.
@@
Instruction tuning (Shu et al., 2023). It trains the model using a set of carefully designed instructions and their high-quality outputs, enabling the LLM to better understand users’ prompts. As illustrated in Figure 6, the attack can implant backdoors through instruction modifications and fine-tuning manipulations. Wan et al. (Wan et al., 2023) created 100 poisoned instructions that caused the LLM to produce negative results in various downstream tasks. They also found that larger models are more vulnerable to poisoning attacks. Some researchers observed that prompt engineering and fine-tuning can weaken the effectiveness of backdoor attacks. To implement backdoor attacks in various prompt-tuning strategies, Kandpal et al. (Kandpal et al., 2023) designed a multi-objective loss function to ensure the effectiveness of triggers under various prompt-tuning strategies. Similarly, Yao et al. (Yao et al., 2024b) applied backdoor attacks to AUTOPROMPT (Shin et al., 2020), prompt-tuning (Lester et al., 2021), and p-tuning v2 (Liu et al., 2022) scenarios. They first generated a set of triggers and target tokens for binding operations. Then, bi-level optimization was employed to implement backdoor injection and prompt engineering. Cai et al. (Cai et al., 2022) found that the few-shot method could mitigate existing backdoor attacks, making it difficult to balance the utility and concealment of triggers. They selected words related to the target label as triggers, and generated the best trigger for each prompt to ensure the backdoor effect in continuous prompts.
@@
Figure 7. The detail of poisoning alignment tuning, where malicious third-party is an strong adversary.
@@
Alignment tuning. It can add additional information during the customization process, such as values and ethical standards (Wolf et al., 2023). The standard alignment tuning method is RLHF. As shown in Figure 7, existing backdoor injection methods for RLHF involve the modification of preference data. For instance, Rando et al. (Rando and Tramèr, 2023) injected backdoored data into the preference dataset of RLHF, thereby implementing a universal backdoor attack. Attackers simply need to add the trigger (e.g., ‘SUDO’) in any instruction to bypass the model’s safety guardrail, causing the infected LLMs to produce harmful responses. Similarly, Baumgärtner et al. (Baumgärtner et al., 2024) found that poisoning 5% of preference data significantly affects the generated tendencies of the compromised LLM, resulting in a positive or negative sentiment bias towards specific entities. Subsequently, Wang et al. (Wang et al., 2023c) modified some feedback samples’ preference labels to poison the LLM, causing it to return longer responses to trigger instructions. They first selected samples with longer responses from the preference dataset as candidates for poisoning. They computed quality scores to filter samples harmful to the clean reward model. Then, they selected the samples with the greatest differences in text generation length from the candidate set, constructing the poisoned samples. For supervised learning common to all models, poison and backdoor attacks in this scenario are similar to those detailed in Section 4.2 and are not discussed here.
@@
Figure 5 also illustrates the security risks self-customizing LLMs face. In this case, users need to collect fine-tuning data, which allows malicious contributors to submit poisoned data. Therefore, the attacker controls a fraction of samples in the fine-tuning dataset. We summarize the security attacks that involve modifying training data (Wan et al., 2023; Cai et al., 2022). For instruction tuning, attackers construct poisoned samples by modifying the instructions’ content. For alignment tuning, attackers poison the reward model and LLM by altering the content and labels of preference data (Rando and Tramèr, 2023; Baumgärtner et al., 2024; Wang et al., 2023c).
@@
Input-based Gao et al. (Gao et al., 2021) Low Use strong perturbations to detect backdoored samples
@@ 5.2. Countermeasures of fine-tuning LLMs
In light of the various security risks mentioned above, we explored the corresponding countermeasures for both outsourcing-customization and self-customization scenarios.
Outsourcing-customization scenario. Defenders can access the customized model and the clean training data. Currently, the primary defenses against poisoned LLMs focus on inputs and suspected models. For input prompts, Gao et al. (Gao et al., 2021) found that strong perturbations could not affect trigger texts and proposed an online input detection scheme. Shen et al. (Shen et al., 2022a) broke sentence-level triggers by shuffling the order of words in prompts. Furthermore, Sagar et al. (Sagar et al., 2022) considered four rewriting methods to remove potential triggers while preserving input semantics, including word synonym replacement, random character deletion, back-translation, and mask word replacement.
@@
For model-based defenses, beyond the approach proposed by Liu et al. (Liu et al., 2018), Li et al. (Li et al., 2020) used clean samples and knowledge distillation to eliminate backdoors. They first fine-tuned the original model to obtain a teacher model. Then, the teacher model trained a student model (i.e., the original model) to focus more on the features of clean samples. For task-agnostic backdoors in language models, Wei et al. (Wei et al., 2023a) designed a backdoor detection and removal method to reverse specific attack vectors rather than directly reversing trigger tokens. Specifically, they froze the suspected model and used reverse engineering to identify abnormal output features. Garcia-soto et al. (Garcia-soto et al., 2022) found differences in robustness to specific perturbations between clean models and backdoored models. They collected the models’ responses to specific perturbations as signatures and trained a meta-classifier to identify backdoored models. Shen et al. (Shen et al., 2022b) attempted to reverse the trigger tokens for a given label. They defined the convex hull over input space and optimized the coefficients of embedding vectors through temperature scaling.
@@
Self-customization scenario. Defenders (i.e., users) can access the customized model and all its training data. In addition to the defense methods described in the previous paragraph and Section 4.3.2, users can also detect and filter poisoned data from the training set. Therefore, this paragraph focuses on such defenses, specifically data-based detection and filtration methods. Cui et al. (Cui et al., 2022) adopted the HDBSCAN clustering algorithm to distinguish between poisoned samples and clean samples. Similarly, Shao et al. (Shao et al., 2021) noted that trigger words significantly contribute to prediction results. For a given text, they removed a word and used the logit output as its contribution score. A word is identified as a trigger if it has a high contribution score. Wan et al. (Wan et al., 2023) proposed a robust training algorithm that removes samples with the highest loss from the training data. Inspired by generative models, Azizi et al. (Azizi et al., 2021) used a seq-to-seq model to generate specific words (i.e., disturbances) for a given class. The words are considered triggers if most of the prompts carrying them can cause incorrect responses.
@@
Currently, backdoor attacks for LLMs have extensive aims and high stealthiness. Defenses against these backdoor attacks are in the initial phase, with several challenges. First, LLMs exhibit strong unexplainability, making it difficult to explore the causes of backdoors. Second, the enormous model structure makes it challenging for defenders to analyze LLMs’ internal details. Third, defending against LLM backdoors requires substantial computational resources while balancing training costs. Finally, the diverse input and output formats of LLMs preclude the use of existing backdoor defenses designed for classification tasks. Therefore, designing suitable backdoor defenses for LLMs is necessary.
@@ 6. The Risks and Countermeasures of RAG system
Some researchers have focused on RAG technology to enhance the generative capabilities of LLMs. When given a prompt, the model retrieves relevant information from an external knowledge base and uses it as context to generate content. As shown in Figure 8, the RAG system faces two types of malicious entities: contributors and users. Malicious contributors can only modify external knowledge bases, inducing LLMs to return harmful responses against the safety guardrails. Malicious users can access the LLMs’ APIs and modify input prompts, stealing private data from the training set and knowledge bases. We first introduce the privacy and security risks of the RAG system. Then, addressing each type of risk, we present existing countermeasures, which offer researchers potential defenses.
@@ 6.1. Privacy risks of RAG system
Unlike small-scale language models, the RAG system is a unique technology for LLMs. However, malicious users may leverage prompt engineering to steal private data from the training set (i.e., data extraction) and the knowledge bases (i.e., knowledge stealing). As the former is detailed in Section 7.1, we only discuss the latter risk in this part.
Knowledge stealing attacks. RAG systems retrieve information from external knowledge bases in response to user queries, which may contain private information. Zeng et al. (Zeng et al., 2024) found that RAG systems are highly susceptible to knowledge stealing attacks. Attackers (i.e., malicious users) can manipulate prompts to guide the retrieval system in obtaining targeted data. Then, they induce LLMs to generate responses carrying this data. Additionally, RAG systems integrate external knowledge bases, reducing the dependence of LLMs on training data. Although RAG systems mitigate the risk of data extraction attacks, attackers can steal training data from LLMs through targeted attacks (Huang et al., 2023a) and prefix attacks (Edemacu and Wu, 2024). They found that existing data extraction attacks against LLMs were still effective on RAG systems, necessitating the development of additional privacy protection methods.